Jorge Company bottles and distributes B-Lite, a diet soft drink. The beverage is sold for 60 cents per 16-ounce bottle to retailers, who charge customers 90 cents per bottle. For the year 2017, management estimates the following revenues and costs.

Sales $2,220,000 Selling expenses—variable $60,000
Direct materials 470,000 Selling expenses—fixed 50,000
Direct labor 300,000 Administrative expenses—variable 62,000
Manufacturing overhead—variable 440,000 Administrative expenses—fixed 50,000
Manufacturing overhead—fixed 579,200

Compute the contribution margin ratio and the margin of safety ratio. (Round variable cost per bottle to 3 decimal places, e.g. 0.25 and final answers to 0 decimal places, e.g. 25%.)

Contribution margin ratio

enter percentages rounded to 0 decimal places

%
Margin of safety ratio

enter percentages rounded to 0 decimal places

%
Answer #1

Number of units sold = Total sales/Selling price per unit

= $2,220,000/0.60

= 3,700,000

Total variable costs = Direct materials + Direct labor + Variable Manufacturing overhead + Variable Administrative expenses + Variable Selling expenses

= 470,000 + 300,000 + 440,000 + 62,000 + 60,000

= $1,332,000

Variable cost per unit = Total variable cost/Number of units sold

= 1,332,000/3,700,000

= $0.36

Contribution margin per unit = Selling price per unit – Variable cost per unit

= 0.60 - 0.36

= $0.24

Contribution margin ratio = Contribution margin per unit/Selling price per unit

= 0.24/0.60

= 40%

Total fixed costs = Fixed Manufacturing overhead + Fixed Administrative expenses + Fixed Selling expenses

= 579,200 + 50,000 + 50,000

= $679,200

Break even sales in dollars = Fixed cost/Contribution margin ratio

= 679,200/40%

= $1,698,000

Margin of safety in dollar sales = Present sales - Break even sales

= 2,220,000 - 1,698,000

= $522,000

Margin of safety ratio = Margin of safety/Present sales

= 522,000/2,220,000

= 24%

Please ask if you have any query related to the question. Thank you